# Towards Interpretable Deep Generative Models via Causal Representation Learning

### v1

@@ 4.2 Latent causal graphs
In the linear case (also known as a linear structural equation model, or SEM), the model for $p_{\bm{z}}(\bm{z}_{i})$ becomes
$\bm{z}_{i}=A^{T}\bm{z}_{i}+\varepsilon,\quad A=(a_{ij})\in\mathbb{R}^{K\times K},\quad\varepsilon_{j}\perp\!\!\!\perp\operatorname{pa}_{\mathsf{G}}(j).$ (4.2)
Then we can associate to $\bm{z}_{i}$ a latent causal graph $\mathsf{G}$ by reading off the nonzero entries of the matrix $A$: Simply define $\mathsf{G}=(\mathsf{V},\mathsf{E})$ by
[…]
See Figure 5 for an example SEM. It turns out that as long as $\mathsf{G}$ is acyclic, then $\mathsf{G}$ is a Bayesian network (i.e. a DAG) for $\bm{z}_{i}$. In other words, the matrix $A$ explicitly encodes a Bayesian network for $\bm{z}_{i}$.
4.3 Interventions and causality
The edges in a DAG can be interpreted as indicating the causal dependencies in a system: The edge $\bm{z}_{ij}\to\bm{z}_{ik}$ indicates that $\bm{z}_{ik}$ depends causally on $\bm{z}_{ij}$ in the sense that changes to $\bm{z}_{ij}$ propagate to $\bm{z}_{ik}$. Returning to the linear SEM example in (4.2), such causal dependencies can be seen by writing the model component-wise:
$\bm{z}_{ik}=\sum_{j=1}^{K}a_{jk}\bm{z}_{ij}+\varepsilon_{k}.$ (4.4)
Explicitly, $\bm{z}_{ik}$ is a noisy function of its parents.
Interventions and independent mechanisms.
This causal interpretation allows us to reason about how the system changes when we externally manipulate the values of certain variables, i.e. an intervention. An intervention occurs when we fix the value of a certain variable (or a set of variables) to a particular value. When this happens, the causal parents of the intervened variable are ignored, akin to deleting any incoming edges to this node.
To make this concrete, consider again the linear SEM in Figure 5. In equilibrium (in particular, without any external interventions) the value of $\bm{z}_{i3}$ is determined by the values of $\bm{z}_{i1}$ and $\bm{z}_{i2}$. The value of $\bm{z}_{i2}$, in turn, is determined by $\bm{z}_{i1}$, whereas $\bm{z}_{i1}$ is an independent exogenous variable that is not determined by any other $\bm{z}_{ik}$. Now suppose we intervene on $\bm{z}_{i2}$: Then the edge from $\bm{z}_{i1}\to\bm{z}_{i2}$ is broken, that is, the value of $\bm{z}_{i1}$ is no longer relevant for determining $\bm{z}_{i2}$: The value of $\bm{z}_{i2}$ is fixed by external manipulation. This then propagates down to $\bm{z}_{i3}$, whose value is determined by the (unintervened) value of $\bm{z}_{i1}$ and the intervened value of $\bm{z}_{i2}$. The result is a new linear SEM
$\bm{z}_{i1}$ $=\varepsilon_{1}$
$\bm{z}_{i2}$ $=\bm{z}_{i2}^{*}$
$\bm{z}_{i3}$ $=-3\bm{z}_{i1}+3\bm{z}_{i2}^{*}+\varepsilon_{3}$
$\bm{z}_{i4}$ $=\bm{z}_{i3}+\varepsilon_{4},$
where $\bm{z}_{i2}^{*}$ is the value that $\bm{z}_{i2}$ is manipulated to. To denote the new distribution of $\bm{z}_{i3}$ given this intervention on $\bm{z}_{i2}$, the convention is to use do-notation:
$P(\bm{z}_{i3}|\bm{z}_{i1},\mathrm{do}(\bm{z}_{i2}=z_{2}^{*})).$
Implicit in this formalization of interventions is the principle of modularity or independent mechanisms: Intervening on a node changes the local conditional distribution of that node, but leaves the remaining conditional distributions unchanged. See Peters et al., (2017) (Chapter 2) for more discussion on the principle of independent mechanisms.
Latent interventions.
The factor model (Definition 1) makes explicit that intervening on the observed features $\bm{x}_{i}$ is not meaningful. Instead, we are interested in intervening on latent variables, which might seem strange at first glance. After all, we do not directly observe latent variables. In many applications, however, this is quite natural: For example, recall the motivating examples in Section 2.
When the observed features are pixels (Figure 1(b)), intervening on individual pixels is not meaningful outside of toy applications. In practice, however, we can obtain additional images or data augmentations in which the object of the image is translated or rotated, or the background is changed. Each of these changes can be viewed as an intervention. However, it is difficult to model such interventions at the pixel level; e.g. in different images, different subsets of pixel values correspond to the background. A solution to this problem is to learn a latent representation $\bm{z}_{i}$ of each image $\bm{x}_{i}$ where the dimensions $\bm{z}_{ik}$ have a consistent interpretation for all $i$; e.g. $\bm{z}_{ik}$ corresponds to the rotation of the images. This type of data is readily available in applications, and can even be easily generated with the aid of simulators (e.g. Von Kügelgen et al.,, 2021; Brehmer et al.,, 2022).
In genomics, the observations are gene expression levels (Figure 1(a)). Genes may coordinate as part of a biological pathway; collectively, these genes may induce effects in other variables. As such, it may be more efficient to aggregate these gene subsets into single causal variables representing gene programs or biological processes. Other examples of latent interventions include medical treatments (drugs, surgeries), radiation exposure, or other epigenetic factors (DNA methylation, histone modification, environmental factors).
We refer to an intervention where a single dimension of $\bm{z}$ is intervened upon as a single-node intervention. For a linear SEM, we formalize this as follows.
Definition 2 (Single-node latent interventions).
A interventional distribution $p_{\bm{z}^{(e)}}$ is a single node intervention of the latent distribution $p_{\bm{z}}$ if there is a single target node $t(e)\in[K]$, and the assigning equation of $z^{(e)}_{i,t(e)}$ is:
$z_{i,t(e)}^{(e)}$ $=(A^{(e)}\bm{z}_{i}^{(e)})_{t(e)}+\varepsilon_{i,t(e)}+c^{(e)}$ (4.5)
while leaving all other equations $\bm{z}_{i,k}$ for $k\neq t(e)$ unchanged. We assume $a_{t(e),k}^{(e)}\neq 0$ only if $k\to t_{e}$ in $\mathsf{G}$; that is, no parents are added. The variable $c^{(e)}$ denotes a shift of the mean; the variance of the noise $\varepsilon_{i,t(e)}$ may also be different to that of $p_{\bm{z}}$. An intervention is called perfect if $a_{t(e),k}^{(e)}=0$ for all $k$ i.e. all connections to former parents are removed.
Intervention targets can be either known or unknown: When intervention targets are known, there is no ambiguity about the mapping $t(e)$ defined above. A considerably more challenging setting arises when intervention targets are unknown, meaning $t(e)$ is not known. In the setting of CRL, since the intervention targets are latent and unobserved, these interventions must also be unknown out of necessity—in fact, we may not even know the correct number of latent factors to begin with.
This leads to additional complications in CRL: Although there is a growing literature on learning under unknown interventions (Eaton and Murphy,, 2007; Kocaoglu et al.,, 2019; Jaber et al.,, 2020; Squires et al.,, 2020; Castelletti and Peluso,, 2022; Hägele et al.,, 2022; Faria et al.,, 2022), this line of work focuses on interventions on measured variables (i.e. $\bm{x}_{ij}$). When interventions are both unknown and latent (i.e. $\bm{z}_{ij}$), much less is known, and this is a developing area of research driven by problems in CRL; e.g. Jiang and Aragam, (2023).
5 What is causal representation learning?
With these preliminaries on factor models and causality out of the way, we are ready to discuss CRL in more precise terms.
Simply put, CRL takes a modern view on factor analysis, driven by the success of generative AI and guided by three core principles as outlined in the introduction: Flexibility, interpretability, and transferability.
  • Flexibility is achieved by allowing (a) correlations between latents, and (b) allowing for general nonlinear relationships through arbitrarily complex neural networks and deep learning;
  • Transferability is achieved through the use of causal models, which are widely believed to yield stable, invariant, transferable predictions across changing contexts and environments.
To capture these three core principles, CRL begins with a general nonlinear factor model
$\bm{x}_{i}=f(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}p(\bm{\varepsilon}),\quad i=1,\dots,n,$ (5.1)
where the latent $\bm{z}_{i}$ are related by some latent causal graph $\mathsf{G}$ and the nonlinearity $f$ is typically estimated with deep neural networks, as in a deep generative model, and causal relationships are inferred through the use of data arising from multiple environments (e.g. across different experimental interventions, time points, etc.). Examples of generative models explicitly of the form (5.1) include diffusion models, variational autoencoders, and flows. Moreover, this class of models continues to dominate leaderboards tracking progress in generative AI (HuggingFace,, 2024; Liu et al.,, 2024).
Remark 1.
The model (5.1) is an assumption on the distribution that can be made independently of whether neural networks are used to fit the data or not; in particular it is not necessary to assume that $f$ itself is a neural network. Of course, in practice we are interested in using neural networks to fit this model nonparametrically.
It is worth contrasting (5.1) with the earlier linear factor model (3.1): The only difference is replacing the linear map $\bm{B}\bm{z}_{i}$ with a nonlinear map $f(\bm{z}_{i})$. Clearly, the former is a special case of the latter. Since even linear factor models are not identifiable, it is clear that (5.1) is also not identifiable. One of the key endeavours in CRL is to determine realizable assumptions under which this model becomes identifiable and estimable.
5.1 Problem definition
To discuss statistical issues such as identifiability, we need to specify the CRL problem a bit more concretely. Given a function class $\mathcal{F}$ and prior family $\mathcal{P}$, we assume the observed data $\bm{x}_{i}\in\mathcal{X}$ is generated from the latent $\bm{z}_{i}\in\mathcal{Z}$ via an unknown mixing function $f\in\mathcal{F}$:
$\bm{x}_{i}=f(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}p(\bm{\varepsilon}),\quad i=1,\dots,n,$ (5.2)
where the latent $\bm{z}_{i}\sim p_{\bm{z}}\in\mathcal{P}$ are generated according to a structural equation model:
$\bm{z}_{i}=A\bm{z}_{i}+D^{1/2}\bm{\nu}_{i},\quad\nu_{ik}\perp\!\!\!\perp\operatorname{pa}(z_{ik}),$ (5.3)
where $A$ is the SEM matrix for some DAG $\mathsf{G}$ (cf. (4.2)) and $D$ is a diagonal matrix with positive entries. This prior on $\bm{z}_{i}$ is denoted by $p_{\bm{z}}\in\mathcal{P}$.
Here, $\mathcal{F}$ is a class of transformations and $\mathcal{P}$ is a family of distributions over $\bm{z}$. For example, in the linear factor analysis model $\mathcal{F}$ is a class of linear maps and $\mathcal{P}$ is typically taken to be the family of Gaussian distributions. The model can then be formalized as the collection of distributions induced by these families, and more general models can be considered by simply expanding $\mathcal{F}$ to include nonlinear transformations or $\mathcal{P}$ to include non-Gaussian distributions. A common choice for $\mathcal{F}$ is the family of invertible and differentiable functions; see Kivva et al., (2022) for a detailed discussion of the implications of various functional restrictions.
The goal is to infer the unknown function $f$ and the prior on the latents $p_{\bm{z}}$ including $A$, which encodes the causal graph of the latent variables. The parameters $(f,p_{\bm{z}},A)$ are important for sampling from the model and understanding the effect of interventions on the observed data, with $A$ being important to understand the causal mechanisms and dependence structure.
Remark 2.
In this review, we focus on models where the latent variables satisfy a linear structural equation model (5.3). Of course, one can consider more general models for $p_{\bm{z}}$ (Jiang and Aragam,, 2023; Varici et al.,, 2023, 2024; von Kügelgen et al.,, 2023); see the end of Section 6 for a brief overview.
Remark 3.
The model (5.1) is a nonlinear factor model, which has a long history in both statistics and ML; see Amemiya and Yalcin, (2001) for a review. See also Zhou et al., (2024); Xu et al., (2023) for recent developments. Although the mathematical models share similarities, what differentiates CRL from prior work on nonlinear factor analysis is the focus on deep learning and causality that drives much of CRL. For example, a key distinction arises from the latent SEM (5.3) and the causal, interventionist semantics it introduces.
5.2 Generation vs. understanding
As noted in the introduction, in applications of generative AI there is a crucial distinction between the problems of generation and understanding that parallels the two cultures of statistical modeling. Since the vast majority of the literature on generative AI focuses on generation, we pause for a moment to outline in more detail these two goals:
  1. 1.
    Generation (density estimation, sampling). Here, we evaluate a model based on its ability to reconstruct the true data density of $\bm{x}_{i}$, and ignore what latents are learned or how they can be interpreted. More formally, generation quality is measured by distributional metrics such as total variation, Wasserstein distance, or Kullback-Liebler (KL) divergence. For example, if $p$ represents the density of $\bm{x}_{i}$, then we can evaluate the quality of a model $\widehat{p}$ with the KL divergence $\textup{KL}(p\,\|\,\widehat{p})$. Closely related to this goal is sampling from $p$, which is a key objective in many applications.
@@ 5.1 Problem definition
    Understanding (parameter estimation, learning). Here, we evaluate a model based on the quality of the latents themselves, which are captured by the model parameters $\bm{z}_{i}$ and $f$. This could mean recovering the precise latent codes $\bm{z}_{i}$ for each sample $\bm{x}_{i}$, the latent causal graph $\mathsf{G}$, or even just the distribution over $\bm{z}_{i}$. In other words, our model should capture meaningful semantics about the generative process.
The distinction between these two goals has two important statistical implications: 1) Generation is significantly simpler than understanding, and can be achieved without understanding; and 2) To understand models, parameters should be identifiable so that they can be consistently estimated as in traditional statistical models.
@@ 5.3 Identifiability
It is clear that (5.1) is not identifiable: Different parameters $(f,p_{\bm{z}})$ lead to the same distribution $p(\bm{x}_{i})$ over the observables. For generation, identifiability is not necessary, however, as we have discussed, we want to go beyond generation to deeper questions of understanding and interpretability. Statistically, we would like to discuss parameter estimation in (5.1), analogous to how factors are estimated in classical linear factor models. While standard in the statistics literature, the relevance of identifiability for ML/AI applications has also been articulated in recent work (D’Amour et al.,, 2020; Reizinger et al.,, 2024).
@@
    Causal identifiability of the causal graph $\mathsf{G}$ and its parameters $A$.
@@
Statistical identifiability can be understood in analogy with factor analysis: This is identifiability of the latent factors themselves. Causal identifiability refers to the distinct problem of identifying the causal relationships between the latent factors as encoded by the latent causal graph $\mathsf{G}$. Of course, both are desirable in practical applications.
@@
These notions of identifiability are distinct: Statistical identifiability is insufficient to imply causal identifiability, and vice versa. That $\mathsf{G}$ cannot be identified from $p_{\bm{z}}$ is a well-known result from graphical models: $\mathsf{G}$ can only be identified up to Markov equivalence. For example, consider two variables $(\bm{z}_{i1},\bm{z}_{i2})$ with joint distribution $p_{\bm{z}}(\bm{z}_{i1},\bm{z}_{i2})$. By the chain rule of probability, $p_{\bm{z}}$ admits two factorizations:
$p_{\bm{z}}(\bm{z}_{i1})p_{\bm{z}}(\bm{z}_{i2}\,|\,\bm{z}_{i1})\quad\text{and}\quad p_{\bm{z}}(\bm{z}_{i2})p_{\bm{z}}(\bm{z}_{i1}\,|\,\bm{z}_{i2}).$
@@
That is, $p_{\bm{z}}$ factorizes with respect to both the graph $\bm{z}_{i1}\to\bm{z}_{i2}$ and $\bm{z}_{i2}\to\bm{z}_{i1}$. Consequently, $\mathsf{G}$ is not identifiable from $p_{\bm{z}}$. The converse that $p_{\bm{z}}$ cannot be identified from $\mathsf{G}$ is a classical result; see Drton et al., (2023) for recent developments.
@@ 6 Identifiability in causal representation learning
In this section, we consider in detail the identifiability of the model (5.2-5.3). First, we review approaches to statistical identification of the model parameters $(f,p_{\bm{z}})$. Then, we review approaches to causal identification, i.e. learning the causal graph between the latent variables $Z$ using interventions.
@@ 6.1 Statistical identifiability: Nonlinear representation learning
In Section 3, we reviewed the identification problem in linear factor analysis. Of course, identification becomes harder if we relax linearity: Informally, we can always re-write (5.2):
[…]
Thus, without restrictions on $\mathcal{F}$ or $\mathcal{P}$, the model is trivially nonidentifiable. Once we impose restrictions on the parameter spaces $\mathcal{F}$ and $\mathcal{P}$, however, identifiability becomes trickier.
To clarify this, let $m_{\#}p_{\bm{z}}$ denote the push-forward measure of $p_{\bm{z}}$ by $m$, i.e. the distribution on $\bm{z}_{i}$ induced by $m$. Then if (i) $f\circ m^{-1}\in\mathcal{F}$ and (ii) $m_{\#}p_{\bm{z}}\in\mathcal{P}$ for some non-identity $m$, then the model is not identifiable. Following the terminology of Xi and Bloem-Reddy, (2023), we refer to such transformations $m$ that satisfy both (i) and (ii) as indeterminancy transformations; the set of indeterminancy transformations is the indeterminancy set.
@@
[…]
It is clear that the identity transformation $m=\mathrm{id}$ is always in the indeterminacy set $\mathcal{A}(\mathcal{F},\mathcal{P})$. When $\mathcal{A}(\mathcal{F},\mathcal{P})$ contains additional transformations, then the model parameters are not identifiable. To resolve identifiability issues, we can again further restrict the function class $\mathcal{F}$, the prior distributions $\mathcal{P}$, or both. We state this more precisely in Definition 4; for a formal measure-theoretic treatment of this heuristic, see Xi and Bloem-Reddy, (2023).
@@
[…]
@@
    For the model parameterized by $(f,p_{\bm{z}})\in\mathcal{F}\times\mathcal{P}$, we say that $f$ is identifiable up to $\mathcal{A}(\mathcal{F},\mathcal{P})$.
@@
    In particular, the model parameters $(f,p_{\bm{z}})$ are strongly identifiable if $\mathcal{A}(\mathcal{F},\mathcal{P})=\{\mathrm{id}\}$; that is, the identity function on $\mathcal{Z}$ is the only indeterminancy transformation.
@@
In many applications, we may relax the requirement of strong identifiability. For example, we are often not interested in the ordering of the latent variables and so the indeterminancy set can contain permutations of the labels (“label switching”). Other common indeterminacies include scaling (diagonal matrices), translations, and orthogonal transformations.
A common goal in machine learning is “disentanglement” (Bengio,, 2013), which roughly means to separate distinct latent factors of variation. Definition 5 below discusses the relationship between identifiability and “disentangling” latent factors of variation.
@@
[…]
@@
where $\sigma$ is a permutation and $g_{k}$ are invertible scalar functions (Ahuja et al.,, 2022; Yao et al., 2024a, ). This is equivalent to the indeterminancy set $\mathcal{A}(\mathcal{F},\mathcal{P})$ containing only element-wise transformations and permutations.
@@
In the earlier ML literature, “disentanglement” referred to finding independent factors of variation (e.g. Bengio,, 2013). Unfortuantely independence is not sufficient for identification, however, a fact that is well-known in the ICA literature (Hyvärinen and Pajunen,, 1999). This was emphasized in the disentanglement context by Locatello et al., (2019). Consequently, the field moved towards identification-based definitions of disentanglement. Eventually, dependent factors of variation began to be considered, leading to the development of causal representation learning.

### v2

@@ 5.1 Problem definition
    Understanding (parameter estimation). Here, we evaluate a model based on the quality of the latents themselves, which are captured by the model parameters $\bm{z}_{i}$ and $f$. This could mean recovering the precise latent variables $\bm{z}_{i}$ for each sample $\bm{x}_{i}$, the latent causal graph $\mathsf{G}$, or even just the distribution over $\bm{z}_{i}$. In other words, our model should capture meaningful semantics about the generative process.
The distinction between these two goals has two important statistical implications: 1) Generation has progressed remarkably without understanding the underlying models, suggesting generation is more tractable than understanding; and 2) To understand models, parameters should be identifiable so that they can be consistently estimated as in traditional statistical models.
@@ 5.3 Identifiability
It is clear that (5.1) is not identifiable: Different parameters $(f,p_{\bm{z}})$ lead to the same distribution $p(\bm{x}_{i})$ over the observables. For generation, identifiability is not necessary, however, as we have discussed, we want to go beyond generation to deeper questions of understanding and interpretability. Statistically, we would like to discuss parameter estimation in (5.1), analogous to how factors are estimated in classical linear factor models. While standard in the statistics literature, the relevance of identifiability for ML/AI applications has also been articulated in recent work (D’Amour et al., 2020; Reizinger et al., 2024).
@@
    Causal identifiability of the causal graph $\mathsf{G}$ and its parameters (e.g. $\bm{A}$).
@@
Statistical identifiability can be understood in analogy with factor analysis; i.e. identifiability of the latent factors. Causal identifiability is the distinct problem of identifying the causal relationships between the latent factors as encoded by the latent causal graph $\mathsf{G}$.
@@
These notions of identifiability are distinct: Statistical identifiability does not imply causal identifiability, and vice versa. That $\mathsf{G}$ cannot be identified from $p_{\bm{z}}$ is a well-known result in graphical models: Simply by the chain rule of probability, $p_{\bm{z}}$ can be factorized over different graphs, each of which encodes (potentially) different conditional independence statements. This leads to the notion of the Markov equivalence class, which collects all the factorizations of $p_{\bm{z}}$ that encode exactly the same conditional independence relations as $p_{\bm{z}}$. For example, consider two dependent variables $(z_{i1},z_{i2})$ with joint distribution $p_{\bm{z}}(z_{i1},z_{i2})$. By the chain rule of probability, $p_{\bm{z}}$ admits two factorizations:
$p_{\bm{z}}(z_{i1})p_{\bm{z}}(z_{i2}\,|\,z_{i1})\quad\text{and}\quad p_{\bm{z}}(z_{i2})p_{\bm{z}}(z_{i1}\,|\,z_{i2}).$
@@
That is, $p_{\bm{z}}$ factorizes with respect to both the graph $z_{i1}\to z_{i2}$ and $z_{i2}\to z_{i1}$. Consequently, $\mathsf{G}$ is not identifiable from $p_{\bm{z}}$ (for more on Markov equivalence, see Lauritzen, 1996; Pearl, 2009). The converse—that $p_{\bm{z}}$ cannot be identified from $\mathsf{G}$—is also classical; see Drton et al. (2023) for recent developments.
In practice, both notions of identifiability are desirable, however, we may be limited to one or the other depending on the available data and assumptions. Whether or not we are interested in causal or statistical identifiability depends on the problem at hand, although in typical applications causal identifiability comes after statistical identifiability; i.e. once we can identify and interpret the representations, we are interested in the effects of manipulating these representations.
5.4 Identifiability vs. interpretability
While identifiability is an important prerequisite for statistical estimation and reliability, it is not sufficient for interpretability. This is well-known: For example, the latent factors recovered by PCA are well-defined but often lack clear interpretation. Here it is helpful to distinguish between two layers of interpretability: First, can these factors be interpreted at all? and second, if they can be interpreted, do they correspond to meaningful concepts? While identifiability partially addresses the first layer, it does not address the second layer.
Moving beyond identifiability, the first layer can also be addressed by leveraging structure in $f$: When $f$ is identifiable, latent factors may be understood by how they relate to the observed features. In linear models, this is reflected in the loadings matrix: If a group of observed features with shared meaning (e.g. genes) load on the same factor, that factor can inherit this interpretation (e.g. the biological function of these genes). Recent work extends this idea to nonlinear settings (Ainsworth et al., 2018; Moran et al., 2022).
The challenge with the second layer is that there is no formal definition of “meaningful”: What is meaningful depends on context and is subjective. For example, size, colour, and shape are nearly universal concepts, however, their precise definitions can vary from one application to the next; e.g. “size” can be interpreted as physical size, file size, or population size. Each of these interpretations measures some aspect of “largeness”, but in incomparable units for different applications.
Thus, to capture the second layer, we often use auxiliary information in the form of concept labels, metadata, or environment annotations that capture this subjective information. This allows us to ask that not only we identify the latent factors and the causal graph, but also that the factors meaningfully capture subjective qualities. This adds an additional constraint on the CRL problem that connects with previous work on interpretable ML (Koh et al., 2020; Schut et al., 2023), with the main difference being the added focus on identifiability in CRL. For example, to identify meaningful concepts, Taeb et al. (2022) require the factors to predict an auxiliary label; Zhang et al. (2023) learn a mapping from intervention label to the latent space; and Rajendran et al. (2024) use concept-conditional environments and contrastive learning to learn interpretable concept representations. These ideas also relate to an emerging literature that explores how to identify human-interpretable concepts in foundation models (such as large language models) that is discussed in more detail in Section 8 (under “Large Language Models”). As this literature is nascent and rapidly developing, we defer a detailed discussion to the papers themselves.
Confounding.  So far we have assumed every latent factor is of interest. In the literature, latent factors have often been interpreted as confounding variables of no inherent interest (e.g. batch effects, Wang and Blei, 2019; Ćevid et al., 2020). It is to be expected that there will be both spurious and meaningful latent factors in data; how to handle and understand these factors is an open question in CRL. Some promising directions in this regard include: allowing both dense and sparse factor loadings (Chandrasekaran et al., 2011; Zhao et al., 2016; Moran et al., 2021), and utilizing interventional data (Luo et al., 2025).
@@ 6 Identifiability in causal representation learning
In this section, we consider in detail the identifiability of the model (5.1-5.2). First, we discuss approaches to statistical identification of the model parameters $(f,p_{\bm{z}})$. Then, we discuss approaches to causal identification, i.e. learning the causal graph between the latent variables using interventions.
@@ 6.1 Statistical identifiability: Nonlinear representation learning
In Section˜3, we reviewed the identification problem in linear factor analysis. Of course, identification becomes harder if we relax linearity: Informally, we can always re-write (5.1):
$\bm{x}_{i}$ $=f(m^{-1}(m(\bm{z}_{i})))+\bm{\varepsilon}_{i},\quad\text{for any invertible $m\mathrel{\mathop{\ordinarycolon}}\mathcal{Z}\to\mathcal{Z}$.}$ (6.1)
Thus, without restrictions on $\mathcal{F}$ or $\mathcal{P}$, the model is trivially nonidentifiable.
To clarify this, let $m_{\#}p_{\bm{z}}$ denote the push-forward measure of $p_{\bm{z}}$ by $m$, i.e. the distribution on $\bm{z}_{i}$ induced by $m$. Then if both $f\circ m^{-1}\in\mathcal{F}$ and $m_{\#}p_{\bm{z}}\in\mathcal{P}$ for some non-identity $m$, then the model is not identifiable. Following the terminology of Xi and Bloem-Reddy (2023), we refer to such transformations $m$ that satisfy both (i) and (ii) as indeterminancy transformations; the set of indeterminancy transformations is the indeterminancy set.
@@
Given $(f,p_{\bm{z}})\in\mathcal{F}\times\mathcal{P}$, $m\mathrel{\mathop{\ordinarycolon}}\mathcal{Z}\to\mathcal{Z}$ is an indeterminancy transformation if $f\circ m^{-1}\in\mathcal{F}$ and $m_{\#}p_{\bm{z}}\in\mathcal{P}$. The indeterminancy set, $\mathcal{A}(\mathcal{F},\mathcal{P})$, is the set of all indeterminancy transformations of $(\mathcal{F},\mathcal{P})$.
It is clear that the identity transformation $m=\mathrm{id}$ is always in the indeterminacy set $\mathcal{A}(\mathcal{F},\mathcal{P})$. When $\mathcal{A}(\mathcal{F},\mathcal{P})$ contains additional transformations, then the model parameters are not identifiable. To resolve identifiability issues, we can again further restrict the function class $\mathcal{F}$, the prior distributions $\mathcal{P}$, or both. We state this more precisely in Definition˜4; for a formal measure-theoretic treatment, see Xi and Bloem-Reddy (2023).
@@
Let $\mathcal{P}$ be a family of probability distributions on $\mathcal{Z}$ and $\mathcal{F}$ be a family of functions $f\mathrel{\mathop{\ordinarycolon}}\mathcal{Z}\rightarrow\mathcal{X}$.
@@
    For the model parameterized by $(f,p_{\bm{z}})\in\mathcal{F}\times\mathcal{P}$, we say that $f$ is identifiable (from $f_{\#}p_{\bm{z}}$) up to $\mathcal{A}(\mathcal{F},\mathcal{P})$.
@@
    The model parameters $(f,p_{\bm{z}})$ are strongly identifiable (from $f_{\#}p_{\bm{z}}$) if $\mathcal{A}(\mathcal{F},\mathcal{P})=\{\mathrm{id}\}$; i.e. the identity function on $\mathcal{Z}$ is the only indeterminancy transformation.
@@
Furthermore, we often relax the requirement of strong identifiability. For example, we may not be interested in the ordering of the latent variables and so the indeterminancy set can contain permutations of the labels (“label switching”). Other common indeterminacies include scaling (diagonal matrices), translations, and orthogonal transformations. A common goal in machine learning is “disentanglement” (Bengio, 2013), which roughly means to separate distinct latent factors of variation. Definition˜5 below helps formalize the relationship between identifiability and “disentangling” latent factors of variation.
@@
Consider a learned representation $\widehat{\bm{z}}=h(\bm{x})$ where $h\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{D}\to\mathbb{R}^{K}$ is an estimate of $f^{-1}$. Then $\widehat{\bm{z}}$ is said to be “disentangled” with respect to the ground-truth factors $\bm{z}$ if $\widehat{\bm{z}}$ is an element-wise transform of $\bm{z}$, up to permutation:
@@
where $\sigma$ is a permutation and $g_{k}$ are invertible scalar functions (Ahuja et al., 2022; Yao et al., 2024a). This is equivalent to the indeterminancy set $\mathcal{A}(\mathcal{F},\mathcal{P})$ containing only element-wise transformations and permutations.
@@
In the earlier ML literature, “disentanglement” referred to finding independent factors of variation (e.g. Bengio, 2013). Unfortunately, independence is insufficient for identification, a fact that is well-known in the ICA literature (Hyvärinen and Pajunen, 1999). This was emphasized in the disentanglement context by Locatello et al. (2019). These results rely on the classical Darmois construction, which uses an orthogonal transformation to keep the prior $p_{\bm{z}}$ invariant while rotating the mixing function $f$. Consequently, there was a shift towards identification-based definitions of disentanglement. Eventually, dependent factors of variation began to be considered alongside the development of CRL.